Help please!
The following problem has two solutions:
7^(x^2)=(7^5x)/(49^3)
FYI,I know that the sum of the two solutions is 5.
I just need a brush-up on some exponential and logarithmic properties so I can know how to do this stuff, thanks!
J

Answer by stanbon(75887)

7^(x^2)=(7^5x)/(49^3)
------------
7^(x^2) = (7^(5x)/(7^2)^3)
---
7^x^2 = 7^(5x)/7^6
---
7^(x^2) = 7^(5x-6)
---
x^2 = 5x-6
x^2-5x+6 = 0
(x-6)(x+1) = 0
---
x = 6 or x = -1
========================
Cheers,
Stan H.
========================